# Evaluating Scoring Bias in LLM-as-a-Judge

### v1

@@ Introduction
(3) Judge models have their own scoring tendencies. Figure 5 presents the scoring tendencies of different judge models under various scoring biases on the same dataset FLASK in the form of heatmaps. It is obvious that each judge model exhibits its own specific scoring preferences. Under unbiased conditions, DeepSeek-V3-671B assigns a score of 5 to more than half of the instruction-response pairs. Compared to the human scoring distribution, Mistral-Small-24B-Instruct has a stronger tendency to score a 4 to the samples. Other judge models show smoother scoring distribution shifts, but certain scores are still preferred. GPT-4o demonstrates a preference for score 4, while Qwen3-32B favors score 5. An examination of Figure 5 reveals that score rubrics order bias and score IDs bias have a certain impact on the scoring tendencies of the judge models; however, they do not fundamentally alter the overall score distribution. An exception is Qwen3-8B, whose score distribution is more significantly influenced by biases, possibly due to its small size. It is particularly evident that the reference answer score bias exerts a substantial impact on the score distribution, which will be analyzed in detail in the following sections. Our experimental results show that judge models inherently exhibit biases in their scoring tendencies. Therefore, debiasing the judge models is crucial to ensuring the fairness and reliability of the evaluation.
Figure 5: Heatmaps of judge models’ score distribution with all bias types on FLASK. The horizontal axis is the score, where -1 represents the failed judgment. The vertical axis is the perturbation method. The top lines of all heatmaps are the human scoring tendencies for comparison, and the second lines are results from unbiased judge models.
Exploratory Experiments and Discussion
@@
For advanced or large size judge models, application of scoring prompt that differ from usual preferences are worth to be considered.
@@
In almost all scoring-based evaluation tasks, Arabic numerals are commonly used as scoring IDs. The score rubrics order typically follows a sequential order from score 1 to score 5. These choices align with human writing conventions and cognitive preferences in the design of scoring prompts. However, our experimental results shown in Table 3 indicate that the scoring prompts designed to align with the usual preferences do not necessarily produce the most accurate scoring results. As mentioned earlier, the performance of the GPT-4o judge model in scoring evaluations improves when Roman numerals are used as scoring IDs. In more than half of the cases, using letter grades as scoring Ids improves the scoring accuracy of DeepSeek-V3-671B. Regarding the score rubrics order bias, adopting a descending order improves the scoring accuracy of GPT-4o, Qwen3-32B, and Mistral-Small-24B-Instruct in most instances. Intuitively, employing a random scoring rubrics order often tends to negatively impact the scoring accuracy of all judge models. Based on the above analysis, we recommend incorporating certain intentional deviations from conventional practices in the design of scoring prompts.
The reference answer score bias perform intriguing impacts on the robustness of scoring-based evaluation and play a critical role.
@@
The impact of reference answer score bias on the scoring accuracy and scoring tendency of the judge model is presented in Figure 4 and Figure 5. The experimental results indicate that, compared to the score rubrics order bias and the score IDs bias, the influence of the reference answer score bias on scoring robustness is both more pronounced and demonstrates stronger regularity. Figure 4 illustrates the impact of introducing reference answers with different scores in the input prompt on the scoring accuracy of the judge model. For GPT-4o, DeepSeek-V3-671B, and Mistral-Small-24B-Instruct, the scoring accuracy in most cases improves when reference answers are introduced. However, for the two Qwen judge models, only reference answers with scores of 4 or 5 contribute to enhancing scoring performance. Figure 4 also reveals an interesting phenomenon: the higher the score of the reference answer, the better the scoring accuracy of the judge model. Although DeepSeek-V3-671B, Qwen3-32B, and Mistral-Small-24B-Instruct do not strictly adhere to this pattern, the correlation between their scores and human scores similarly exhibits an upward trend as the reference answer score increases.
@@
Figure 5 shows that the score of the reference answer has a significant impact on the scoring tendency of the judge model. Specifically, the scoring results from DeepSeek-V3-671B, Qwen3-32B, and Qwen3-8B appear to be highly influenced by the score of the reference answer. When the score of the reference answer falls within the range of 1 to 4, these models tend to assign similar scores to the evaluated instruction-response pairs. Interestingly, when a reference answer with a score of 5 is introduced, the scoring behavior of the judge model appears to become more rational. This is particularly notable for DeepSeek-V3-671B, which exhibits a strong tendency to assign a score of 5 without bias. However, the introduction of a reference answer scored at 5 mitigates this phenomenon.
@@
In summary, the above analysis reveals that the score of the reference answer introduced in the prompt has a significant impact on the judge model’s scoring behavior. In particular, when reference answers with a score of 5 are introduced, the models achieve the highest scoring accuracy without disproportionately influencing their scoring preferences. Therefore, it can be concluded that when incorporating reference answers into the scoring-based judge prompt, a full-marked reference answer is the best choice. Interestingly, this conceptually aligns with the principles of traditional statistical metrics, such as BLEU and ROUGE, which emphasize the critical role of golden answers in evaluation.
@@
Although our work offers notable conclusions and insights, some limitations still remain for future work.
Mitigating scoring bias needs further exploration. Although there are no dedicated and fancy methods for addressing scoring bias, intuitive and straightforward approaches, such as scoring multiple times followed by majority voting or averaging the scores, require empirical validation for their effectiveness in mitigating scoring bias.
Additionally, although our work demonstrates the existence of scoring bias in LLM-as-a-Judge, the underlying causes of scoring bias remain to be validated. Approaches such as training data analysis and information flow observation (Yuan et al. 2025) may help identify the reason scoring bias exists within or outside the model.
@@ Conclusion
In this work, we propose scoring bias in LLM-as-a-Judge, defined as scoring judgments shift when judge models are perturbed. We design an evaluation framework, including three types of scoring bias: score rubrics order bias, score IDs bias, and reference answer score bias. We select four LLM-as-a-Judge benchmarks and expand them as our evaluation datasets through data synthesis. Our evaluation metrics can quantify the impact of scoring bias on the scoring accuracy and scoring tendencies of the judge models. Our experimental results clearly demonstrate that scoring bias can affect the robustness and reliability of existing scoring-based judge models. Judge models with different capabilities, sizes, and architectures exhibit different sensitivities to various scoring biases. Our further exploratory experiments and discussions reveal several intriguing phenomena. In summary, our work provides detailed evidence and valuable recommendations for evaluating and improving the scoring reliability of LLM-as-a-Judge.

### v2

@@ Introduction
(3) Judge models have their own scoring tendencies. Figure 6 presents the scoring tendencies of different judge models under various scoring biases on the FLASK dataset in the form of heatmaps. It is obvious that each judge model exhibits its own specific scoring preferences. Under unbiased conditions, DeepSeek-V3-671B assigns a score of 5 to more than half of the instruction-response pairs. Compared to the human scoring distribution, Mistral-Small-24B-Instruct has a stronger tendency to assign a score of 4 to the samples. Other judge models show smoother scoring distribution shifts, but certain scores are still preferred. GPT-4o demonstrates a preference for a score of 4, while Qwen3-32B favors a score of 5. An examination of Figure 6 reveals that score rubric order bias and score ID bias have a certain impact on the scoring tendencies of the judge models; however, they do not fundamentally alter the overall score distribution. An exception is Qwen3-8B, whose score distribution is more significantly influenced by biases, possibly due to its small size. It is particularly evident that the reference answer score bias exerts a substantial impact on the score distribution, which will be analyzed in detail in the following sections. Our experimental results show that judge models inherently exhibit biases in their scoring tendencies. Therefore, debiasing the judge models is crucial to ensuring the fairness and reliability of the evaluation.
Figure 6: Score distributions of LLM judges under various prompt perturbations on the FLASK dataset. The horizontal axis represents the assigned score (where -1 indicates a failed judgment), and the vertical axis lists the different perturbation methods applied. For reference, the top row in each heatmap (Human) shows the distribution of human scores. The second row (None) shows the model’s score distribution for the original, unperturbed prompt.
Methods for Scoring Bias Mitigation
@@
For advanced or large size judge models, applying scoring prompt that deviate from conventional preferences is worth considering.
@@
In almost all scoring-based evaluation tasks, Arabic numerals are commonly used as scoring IDs. The score rubric order typically follows a sequential order from score 1 to score 5. These choices align with human writing conventions and cognitive preferences in the design of scoring prompts. However, our experimental results shown in Table 3 indicate that the scoring prompts designed to align with the usual preferences do not necessarily produce the most accurate scoring results. As mentioned earlier, the performance of the GPT-4o judge model in scoring evaluations improves when Roman numerals are used as scoring IDs. In more than half of the cases, using letter grades as scoring IDs improves the scoring accuracy of DeepSeek-V3-671B. Regarding the score rubric order bias, adopting a descending order improves the scoring accuracy of GPT-4o, Qwen3-32B, and Mistral-Small-24B-Instruct in most instances. Intuitively, employing a random scoring rubric order tends to negatively impact the scoring accuracy of all judge models. Based on our analysis, we propose a core design principle for scoring prompts: intentional yet meaningful deviation from convention. Effective strategies include introducing structured, unconventional formats, such as presenting scoring rubrics in a descending order or using non-numeric identifiers like letter grades (A, B, C). However, the key is to maintain clarity and logical coherence. Therefore, chaotic perturbations—such as a fully randomized rubric order or non-hierarchical identifiers—should be avoided. Moreover, inspired by previous work (Liu et al. 2024c; Park et al. 2024), perturbations applied to scoring prompt also show potential as debiasing strategies. For example, effective approaches include implementing multiple variations of the scoring rubric order or IDs and then aggregating the results, or adversarially training the models on samples enhanced with scoring bias perturbations.
Using full-mark reference answers can stabilize scoring preferences and achieve optimal accuracy.
@@
The impact of reference answer score bias on the scoring accuracy and scoring tendency of the judge model is presented in Figure 5 and Figure 6. The experimental results indicate that, compared to the score rubric order bias and the score ID bias, the influence of the reference answer score bias on scoring robustness is both more pronounced and more regular. Figure 5 illustrates the impact of introducing reference answers with different scores in the input prompt on the scoring accuracy of the judge model. For GPT-4o, DeepSeek-V3-671B, and Mistral-Small-24B-Instruct, the scoring accuracy in most cases improves when reference answers are introduced. However, for the two Qwen judge models, only reference answers with scores of 4 or 5 contribute to enhancing scoring performance. Figure 5 also reveals an interesting phenomenon: the higher the score of the reference answer, the better the scoring accuracy of the judge model. Although DeepSeek-V3-671B, Qwen3-32B, and Mistral-Small-24B-Instruct do not strictly adhere to this pattern, the correlation between their scores and human scores similarly exhibits an upward trend as the reference answer score increases.
@@
Figure 6 shows that the score of the reference answer has a significant impact on the scoring tendency of the judge model. Specifically, the scoring results from DeepSeek-V3-671B, Qwen3-32B, and Qwen3-8B appear to be highly influenced by the score of the reference answer. When the score of the reference answer is between 1 and 4, these models tend to assign scores similar to that of the reference answer. Interestingly, when a reference answer with a score of 5 is introduced, the scoring behavior of the judge model appears to become more rational. This is particularly notable for DeepSeek-V3-671B, which exhibits a strong tendency to assign a score of 5 without bias. However, the introduction of a reference answer scored at 5 mitigates this phenomenon.
@@
The above analysis reveals that the score of the reference answer introduced in the prompt has a significant impact on the judge model’s scoring behavior. In particular, when reference answers with a score of 5 are introduced, the models achieve the highest scoring accuracy without disproportionately influencing their scoring preferences. Therefore, it can be concluded that when incorporating reference answers into a scoring-based judge prompt, a full-mark reference answer is the best choice. Interestingly, this conceptually aligns with the principles of traditional statistical metrics, such as BLEU and ROUGE, which emphasize the critical role of golden answers in evaluation.
@@
While our work offers notable conclusions and insights, several limitations remain to be addressed in future work.
Firstly, our evaluation framework has room for further expansion. More types of scoring biases remain to be identified. Building more comprehensive benchmarks to evaluate scoring bias in LLM-as-a-Judge is also warranted, as a limited data distribution may lead to an unfair evaluation.
Secondly, mitigating scoring bias needs further exploration. Although there are no dedicated methods for addressing scoring bias, intuitive and straightforward approaches, such as scoring multiple times followed by majority voting or averaging the scores, require empirical validation for their effectiveness in mitigating scoring bias. Moreover, the elaborate design of scoring prompt templates, the construction of debiased scoring-based judgment datasets (Park et al. 2024), and adversarial training are potential approaches to enhancing the scoring robustness of judge models.
Additionally, although our work demonstrates the existence of scoring bias in LLM-as-a-Judge, the underlying causes of scoring bias remain to be validated. Approaches such as training data analysis and information flow observation (Yuan et al. 2025) may help identify the reasons for scoring bias, whether it originates from within the model or from external factors.
@@ Conclusion
This work investigates scoring bias in LLM-as-a-Judge, where scoring judgments shift due to minor perturbations in judge prompts. To systematically evaluate this, we designed a framework that perturbs score rubric order, score IDs, and reference answers, quantifying the bias using scoring accuracy and scoring tendency. Experiments on synthesized datasets expanded from four benchmarks confirm that scoring bias significantly undermines the robustness and reliability of existing LLM judges, with varying sensitivity across different models. To mitigate this, we recommend unconventional prompt designs—such as descending rubric order or using letter or Roman numeral IDs—and providing a full-mark reference answer when applicable, both of which effectively improve performance. In summary, our work provides systematic evidence and actionable recommendations for evaluating and enhancing the reliability of LLM judges.